Given f(x) = x^2 + 3x – 4 and g(x) = x^3 – 2x + 2, find the following.
(f/g)(-2)

Answer by ewatrrr(24785)


Hi
f(x) = x^2 + 3x – 4 , f(-2) = -6
g(x) = x^3 – 2x + 2 , g(-2) = -2
(f/g)(-2) = -6/-2 = 3